# Test-Time Anchoring for Discrete Diffusion Posterior Sampling

### v1

@@ 5 Conclusion
We introduce Anchored Posterior Sampling (APS), a training-free posterior sampler with discrete diffusion foundation models. Grounded in our theoretical bounds, APS enables the reuse of a pretrained denoiser without task-specific retraining. APS is built on two algorithmic innovations: Quantized Expectation, which provides fine-grained, gradient-like guidance in discrete spaces, and Anchored Remasking, which adaptively decodes important tokens early in the reverse process. Together, these yield an efficient, scalable, and training-free posterior sampler.
Extensive experiments on linear and nonlinear inverse problems demonstrate that APS achieves state-of-the-art results among discrete samplers and is competitive with continuous baselines—while operating at significantly lower inference cost. Beyond inverse problems, APS also unlocks new capabilities such as training-free stylization, underscoring the flexibility of discrete diffusion models when paired with effective posterior sampling. We believe this work establishes discrete diffusion as a practical and scalable alternative for posterior sampling, with promising extensions to video, multimodal generation, and other structured domains.
@@ Acknowledgments
Reproducibility Statement
Our experiments are built upon the publicly available MMaDA codebase (Yang et al., 2025). All modifications and implementation details are described in Appendix B.1, which includes the pseudocode in Algorithm 1 and the specific parameters used for every experiment. Furthermore, Appendix B.2 provides comprehensive ablation studies and hyperparameter sweeps (Table 4). The experiments use the widely-used public datasets FFHQ (Karras et al., 2019) and ImageNet (Deng et al., 2009). The combination of a public codebase and datasets, along with our detailed Algorithm 1 and description (§B.2), should ensure that our results are readily reproducible.
Ethics Statement
The proposed method for controlled image editing contributes to the democratization of advanced image editing. While this has positive societal benefits, we acknowledge the dual-use nature of generative models and the potential for misuse or misinterpretation of their outputs.
A primary concern arises in the context of inverse problems such as super resolution, inpainting, and deblurring. Our method generates a plausible, high-quality image, but this output is a sample from a posterior distribution and does not represent a unique or guaranteed reconstruction of the original ground truth. A misunderstanding of this core limitation could lead to dangerous assumptions, particularly in sensitive applications like forensic analysis, where a generated image might be misinterpreted as factual evidence. We therefore emphasize that the method’s intended purpose is to enhance perceptual realism for creative or aesthetic applications, and it is not suitable for tasks requiring high-fidelity reconstruction or person identification.
In stylization tasks, there is a risk of unintentional information leakage, where details from the source content image could persist in the stylized output. This could lead to the inadvertent disclosure of sensitive information that a user did not intend to share.
Appendix A Additional Theoretical Results
@@
This appendix develops complementary theory for masked discrete diffusion posterior sampling. We first derive a pathwise variational bound for training a likelihood-tilted reverse process (Theorem A.1), showing that $\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})$ is upper bounded by a reconstruction term, a sum of token-wise KL-divergence terms, and a sequence of measurement likelihood-based tilting terms. We then specialize this analysis to the training-free setting where the token-to-image decoder is shared between unconditional generation and posterior sampling, yielding a bound for test-time anchored posterior sampling (Theorem A.2). We provide theoretical insights drawn from each theorem in Implication subsections after the corresponding proofs.
Theorem A.1 (Discrete Diffusion Posterior Sampling(DDPS)).
@@
Let $Z_{0:1}\!=\!\{Z_{t(i)}\}_{i=0}^{T}$ with $t(i)=i/T$ and s(i)=(i1)/Ts(i)=(i-1)/T be the latent path of a masked discrete diffusion model, and let $q(Z_{0:1}|{\mathbf{x}})$ be the forward noising law from (3). Consider reverse kernels and a terminal decoder that factorize as follows:
@@
$p_{\varphi}(Z_{s}^{l}|Z_{t},{\mathbf{y}})\;\coloneqq\;q\!\big(Z_{s}^{l}|Z_{t}^{l},{\mathbf{x}}_{\varphi}(Z_{t})\big)\,q\!\big({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t})\big).$
@@
$-\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})$ $\leq{\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)\coloneqq\mathbb{E}_{q(Z_{0:1}|{\mathbf{x}})}[-\log p_{\varphi}({\mathbf{x}}|Z_{0})]$
$\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big].$ (8)
@@
Expanding the tilted reverse kernel using the parameterization $p_{\varphi}(Z_{s(i)}|Z_{t(i)},{\mathbf{y}})\coloneqq q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\,q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))$, we can rewrite the last term as
$\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\Big[\log\frac{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))}-\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\Big].$
@@
$-\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})$ $\leq\underbrace{\mathbb{E}_{q(Z_{0:1}|{\mathbf{x}})}[-\log p_{\varphi}({\mathbf{x}}|Z_{0},{\mathbf{y}})]+\mathrm{KL}\!\big(q(Z_{1}|{\mathbf{x}})\,\|\,p_{\varphi}(Z_{1}|{\mathbf{y}})\big)}_{\text{reconstruction + boundary KL}}$
@@
$\quad-\underbrace{\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big]}_{\text{likelihood tilt terms}}\coloneqq{\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi).$
@@
${\mathcal{L}}_{\text{PNELBO}}({\mathbf{x}},{\mathbf{y}};\varphi)$ $\coloneqq\mathbb{E}_{q(Z_{0:1}|{\mathbf{x}})}\!\big[-\log p_{\varphi}({\mathbf{x}}|Z_{0},{\mathbf{y}})\big]$
@@
The remaining likelihood terms,
$\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big],$
capture the effect of incorporating observations $\mathbf{y}$} into posterior sampling. Thus, the overall objective is the PNELBO plus a sequence of likelihood corrections at every step.
@@
$\mathrm{KL}\!\big(q(Z_{s(i)}\!\mid Z_{t(i)},{\mathbf{x}})\,\|\,q(Z_{s(i)}\!\mid Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big).$
@@
$q(Z_{s(i)}={\mathbf{m}}\mid Z_{t(i)}={\mathbf{m}},{\mathbf{x}})=\frac{1-\alpha_{s(i)}}{1-\alpha_{t(i)}}\!,\qquad q(Z_{s(i)}={\mathbf{x}}\mid Z_{t(i)}={\mathbf{m}},{\mathbf{x}})=\frac{\alpha_{s(i)}-\alpha_{t(i)}}{1-\alpha_{t(i)}}.$
@@
$\mathrm{KL}\!\big($ $q(Z_{s(i)}\!\mid Z_{t(i)}\!=\!{\mathbf{m}},{\mathbf{x}})\,\|\,q(Z_{s(i)}\!\mid Z_{t(i)}\!=\!{\mathbf{m}},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big)$
@@
$\mathrm{KL}\!\big(q(Z_{s(i)}\!\mid Z_{t(i)}\!=\!{\mathbf{m}},{\mathbf{x}})\,\|\,q(Z_{s(i)}\!\mid Z_{t(i)}\!=\!{\mathbf{m}},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big)=\frac{\alpha_{s(i)}-\alpha_{t(i)}}{1-\alpha_{t(i)}}\,\mathrm{CE}({\mathbf{x}},{\mathbf{x}}_{\varphi}(Z_{t(i)})).$
Case II: $Z_{t(i)}\neq{\mathbf{m}}$. When the current token is already unmasked, the posterior is deterministic: $q(Z_{s(i)}\!=\!Z_{t(i)}\mid Z_{t(i)}\neq{\mathbf{m}},{\mathbf{x}})=1$. Thus,
$\mathrm{KL}\!\big(q(Z_{s(i)}\!\mid Z_{t(i)}\neq{\mathbf{m}},{\mathbf{x}})\,\|\,q(Z_{s(i)}\!\mid Z_{t(i)}\neq{\mathbf{m}},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big)=0.$
@@
$\mathrm{KL}\!\big(q(Z_{s(i)}\!\mid Z_{t(i)},{\mathbf{x}})\,\|\,q(Z_{s(i)}\!\mid Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big)=\frac{\alpha_{s(i)}-\alpha_{t(i)}}{1-\alpha_{t(i)}}\;\sum_{l=1}^{L}\mathrm{CE}\!\big({\mathbf{x}}^{l},{\mathbf{x}}^{l}_{\varphi}(Z_{t(i)})\big)\bm{1}_{\{Z_{t(i)}^{l}={\mathbf{m}}\}}.$
Combining this with the likelihood tilt terms, we get
${\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ $=\mathbb{E}_{q(Z_{0:1}|{\mathbf{x}})}[-\log p_{\varphi}({\mathbf{x}}|Z_{0},{\mathbf{y}})]+\mathrm{KL}\!\big(q(Z_{1}|{\mathbf{x}})\,\|\,p_{\varphi}(Z_{1}|{\mathbf{y}})\big)$
@@
$\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big].$
@@
${\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ $=\mathbb{E}_{q(Z_{0:1}|{\mathbf{x}})}[-\log p_{\varphi}({\mathbf{x}}|Z_{0},{\mathbf{y}})]$
@@
$\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big].$
@@
${\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ $=\mathbb{E}_{q(Z_{0:1}|{\mathbf{x}})}[-\log p_{\varphi}({\mathbf{x}}|Z_{0})]$
@@
$\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big],$
which completes the proof of the statement. ∎
Implications. Theorem A.1 provides a principled upper bound on the negative log-posterior likelihood in discrete masked diffusion models.
  • Trainable upper bound. The pathwise upper bound ${\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ in (A.1) provides a principled training criterion for discrete diffusion posterior sampling.
@@
  • Data-consistency via tilt. The additive terms $\sum_{i}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))]$ encourage reverse transitions that produce intermediate predictions consistent with the measurement model, integrating task specific information at every denoising step.
@@
  • Compatibility with efficient parameterizations. Because (A.1) is written in terms of token-wise categoricals, it directly supports time-independent or lightweight parameterizations (e.g., shared denoisers), helping scalability to long sequences and high resolution.
Theorem A.2 (Test-time Anchored Posterior Sampling).
@@
${\mathcal{L}}_{\mathrm{APS}}({\mathbf{x}},{\mathbf{y}};\varphi)\coloneqq{\mathcal{L}}_{\mathrm{NELBO}}({\mathbf{x}};\theta)$ $+\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\!\Bigg[\frac{\alpha_{s(i)}-\alpha_{t(i)}}{1-\alpha_{t(i)}}\,\log\frac{\langle{\mathbf{x}}_{\theta}(Z_{t(i)}),{\mathbf{x}}\rangle}{\langle{\mathbf{x}}_{\varphi}(Z_{t(i)}),{\mathbf{x}}\rangle}\,\bm{1}_{\{Z_{t(i)}={\mathbf{m}}\}}\Bigg]$
$-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big].$

### v2

@@ 5 Conclusion
We introduce Anchored Posterior Sampling (APS), a theoretically grounded, training-free sampler that enables the reuse of pretrained discrete diffusion models without task-specific retraining. APS leverages Quantized Expectation to provide gradient-like guidance in discrete spaces and Anchored Remasking to adaptively decode informative tokens early in the denoising process. In our primary vision tasks on linear and nonlinear inverse problems, APS achieves state-of-the-art results among discrete samplers and remains competitive with continuous baselines while using significantly reduced inference cost. Furthermore, APS demonstrates versatility by enabling training-free stylization and steering language models toward better question answering. This work establishes discrete diffusion as a scalable posterior sampling alternative, with promising extensions to multimodal generation and editing.
@@ Acknowledgments
Impact Statement
Our method enables training-free, controlled image editing using pretrained generative models, which can broaden access to advanced creative tools. As with all generative models, it has dual-use potential and may be misapplied if its limitations are misunderstood.
In inverse problems such as super-resolution, inpainting, and deblurring, APS produces a plausible sample from a posterior distribution rather than a unique or guaranteed reconstruction of the original signal. Misinterpreting such outputs as factual reconstructions could be harmful in sensitive settings, including forensic analysis or identity verification. Accordingly, our method is intended for perceptual enhancement tasks, and is not suitable for applications requiring evidentiary reliability.
For stylization, there is a risk that information from the source content image may leak in the generated output, potentially leading to unintended disclosure of sensitive details. Users should therefore exercise caution when applying the method to private or confidential imagery.
Appendix A Additional Related Works
Gaussian pixel-space diffusion methods (Chung et al., 2023; Kawar et al., 2022; Zhu et al., 2023) study inverse problems using diffusion priors trained in the pixel space. These priors are domain-specific, e.g., a model trained on ImageNet must be used for ImageNet tasks, and one trained on FFHQ for FFHQ tasks, resulting in informative but narrow priors. Mixing domains during training convolutes these priors, and at the extreme of internet-scale training, the priors remain valid for generation (Ramesh et al., 2021; Baldridge et al., 2024; Esser et al., 2024; Black Forest Labs, 2024) but become less informative for domain-specific problems. This motivates the challenge of extracting domain-specific priors from general-purpose priors.
Gaussian latent diffusion. PSLD (Rout et al., 2023) introduced posterior sampling with latent diffusion, showing how domain-specific priors can be extracted from general-purpose priors, e.g., large-scale foundation models. This line of work (Chung et al., 2024; Song et al., 2024; Rout et al., 2024, 2025b; Noroozi et al., 2024; Zhang et al., 2025; Chung et al., 2025) uses latent diffusion priors: a single pretrained model can handle multiple domains, enabling inverse problems and semantic edits without retraining, while also being faster and more scalable to high-resolution synthesis. A drawback, however, is that posterior sampling often requires backpropagation through large denoisers (e.g., Flux (Black Forest Labs, 2024), SD3.5 (Esser et al., 2024)), which is prohibitively slow. RB-Modulation (Rout et al., 2025c) alleviates this by framing the problem as stochastic optimal control, directly optimizing the terminal latent state and reducing runtime from minutes (PSLD: $\sim$12 min, P2L: $\sim$30 min, STSL: $\sim$3 min) to under 40 seconds. This efficiency relies on continuous, differentiable latent embeddings, a property that does not extend to discrete diffusion. Addressing this gap motivates the need for new posterior sampling approaches in discrete settings.
Uniform discrete diffusion. Recent works have explored posterior sampling with discrete diffusion. G2D2 (Murata et al., 2024) extends the proximal sampler of RB-Modulation to VQ-diffusion (Gu et al., 2022) using a star-shaped noising process and Gumbel-Softmax dequantization (Gumbel, 1954; Jang et al., 2017; Maddison et al., 2017). While it enables gradient guidance, G2D2 depends on continuous relaxations, requires storing log-probabilities from previous step, and struggles to generalize to purely discrete token embeddings (§4). SGDD (Chu et al., 2025) instead proposes a split Gibbs sampler with Hamming-distance reweighting and rejection sampling via Metropolis–Hastings, but its exponential rejection rate restricts their results to low-resolution tasks.
Masked (absorbing) discrete diffusion. While G2D2 and SGDD can, in principle, be adapted to masked diffusion, they perform poorly with purely discrete token embeddings. In contrast, our method leverages a unified masked discrete diffusion model and introduces two key components: quantized expectation (§3.1) and anchored remasking (§3.2). Together, these yield an efficient and scalable posterior sampler for high-resolution inverse problems. To our knowledge, this is the first inverse problem solver tailored for masked discrete diffusion with purely discrete embeddings, outperforming prior discrete samplers and remaining competitive—often superior—to continuous diffusion methods at substantially lower inference cost (§4).
Appendix B Additional Theoretical Results
@@
This appendix develops complementary theory for masked discrete diffusion posterior sampling. We first derive a pathwise variational bound for training a likelihood-tilted reverse process (Theorem B.1), showing that $\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})$ is upper bounded by a reconstruction term, a sum of token-wise KL-divergence terms, and a sequence of measurement likelihood-based tilting terms. We then specialize this analysis to the training-free setting where the token-to-image decoder is shared between unconditional generation and posterior sampling, yielding a bound for test-time anchored posterior sampling (Theorem B.2). We provide theoretical insights drawn from each theorem in Implication subsections after the corresponding proofs.
Theorem B.1 (Discrete Diffusion Posterior Sampling(DDPS)).
@@
Let $Z_{0:1}\!=\!\{Z_{t(i)}\}_{i=0}^{T}$ with $t(i)=i/T$ and s(i)=(i1)/Ts(i)=(i-1)/T be the latent path of a masked discrete diffusion model, and let $q(Z_{0:1}|{\mathbf{x}})$ be the forward noising law from (3). Consider reverse kernels and a terminal decoder that factorize as
@@
$p_{\varphi}(Z_{s}^{l}|Z_{t},{\mathbf{y}})\;{\propto}\;q\!\big(Z_{s}^{l}|Z_{t}^{l},{\mathbf{x}}_{\varphi}(Z_{t})\big)\,q\!\big({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t}{;Z^{l}_{s}})\big).$
@@
$-\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})$ $\leq{\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)\coloneqq\mathbb{E}_{q(Z_{0}|{\mathbf{x}})}[-\log p_{\varphi}({\mathbf{x}}|Z_{0})]$
$\quad+\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Bigg[\frac{\alpha_{s(i)}-\alpha_{t(i)}}{1-\alpha_{t(i)}}\;\sum_{l=1}^{L}\mathrm{CE}\!\big({\mathbf{x}}^{l},{\mathbf{x}}^{l}_{\varphi}(Z_{t(i)})\big)\bm{1}_{\{Z_{t(i)}^{l}={\mathbf{m}}\}}$
$\qquad\qquad\qquad\qquad-\mathbb{E}_{q({Z_{s(i)}}|Z_{t(i)},{\mathbf{x}})}\Big[\sum_{l=1}^{L}\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}^{l}}))\Big]$
$\qquad\qquad\qquad\qquad{\;+\sum_{l=1}^{L}\log\mathcal{Z}^{l}_{\varphi}(Z_{t(i)},\mathbf{y})\Bigg]}.$ (9)
@@
Expanding the tilted reverse kernel using the parameterization:
$p_{\varphi}(Z_{s(i)}|Z_{t(i)},{\mathbf{y}})$ $\;{=}\;\frac{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\,q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))}{{\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})}},\quad\text{where}$
${\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})}$ ${\;=\;\sum_{{\mathbf{z}}_{s}\in{\mathcal{V}}}q\!\left({\mathbf{z}}_{s}|Z_{t(i)},\mathbf{x}_{\varphi}(Z_{t(i)})\right)\;q\!\left(\mathbf{y}|{\mathbf{x}}_{\varphi}(Z_{t(i)};{\mathbf{z}}_{s})\right),}$
we can rewrite the last term as
$\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\Big[\log\frac{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))}-\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))+{\log\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})}\Big].$
@@
$-\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})$ $\leq\underbrace{\mathbb{E}_{q(Z_{0}|{\mathbf{x}})}[-\log p_{\varphi}({\mathbf{x}}|Z_{0},{\mathbf{y}})]+\mathrm{KL}\!\big(q(Z_{1}|{\mathbf{x}})\,\|\,p_{\varphi}(Z_{1}|{\mathbf{y}})\big)}_{\text{reconstruction + boundary KL}}$
@@
$\quad-\underbrace{\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}{,Z_{s(i)}}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))\big]}_{\text{likelihood tilt terms}}$
${\quad+\underbrace{\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})\big]}_{\text{normalizing factor}}}\coloneqq{\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi).$
@@
${\mathcal{L}}_{\text{PNELBO}}({\mathbf{x}},{\mathbf{y}};\varphi)$ $\coloneqq\mathbb{E}_{q(Z_{0}|{\mathbf{x}})}\!\big[-\log p_{\varphi}({\mathbf{x}}|Z_{0},{\mathbf{y}})\big]$
@@
The likelihood term,
$\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}{,Z_{s(i)}}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))\big],$
capture the effect of incorporating observations $\mathbf{y}$} into posterior sampling. Finally, we have an additive normalizing term that ensures the validity of the reverse transition (6).
@@
$\mathrm{KL}\!\big(q(Z_{s(i)}\!|Z_{t(i)},{\mathbf{x}})\,\|\,q(Z_{s(i)}\!|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big).$
@@
$q(Z_{s(i)}={\mathbf{m}}|Z_{t(i)}={\mathbf{m}},{\mathbf{x}})=\frac{1-\alpha_{s(i)}}{1-\alpha_{t(i)}}\!,\qquad q(Z_{s(i)}={\mathbf{x}}|Z_{t(i)}={\mathbf{m}},{\mathbf{x}})=\frac{\alpha_{s(i)}-\alpha_{t(i)}}{1-\alpha_{t(i)}}.$
@@
$\mathrm{KL}\!\big($ $q(Z_{s(i)}\!|Z_{t(i)}\!=\!{\mathbf{m}},{\mathbf{x}})\,\|\,q(Z_{s(i)}\!|Z_{t(i)}\!=\!{\mathbf{m}},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big)$
@@
$\mathrm{KL}\!\big(q(Z_{s(i)}\!|Z_{t(i)}\!=\!{\mathbf{m}},{\mathbf{x}})\,\|\,q(Z_{s(i)}\!|Z_{t(i)}\!=\!{\mathbf{m}},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big)=\frac{\alpha_{s(i)}-\alpha_{t(i)}}{1-\alpha_{t(i)}}\,\mathrm{CE}({\mathbf{x}},{\mathbf{x}}_{\varphi}(Z_{t(i)})).$
Case II: $Z_{t(i)}\neq{\mathbf{m}}$. When the current token is already unmasked, the posterior is deterministic: $q(Z_{s(i)}\!=\!Z_{t(i)}|Z_{t(i)}\neq{\mathbf{m}},{\mathbf{x}})=1$. Thus,
$\mathrm{KL}\!\big(q(Z_{s(i)}\!|Z_{t(i)}\neq{\mathbf{m}},{\mathbf{x}})\,\|\,q(Z_{s(i)}\!|Z_{t(i)}\neq{\mathbf{m}},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big)=0.$
@@
$\mathrm{KL}\!\big(q(Z_{s(i)}\!|Z_{t(i)},{\mathbf{x}})\,\|\,q(Z_{s(i)}\!|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big)=\frac{\alpha_{s(i)}-\alpha_{t(i)}}{1-\alpha_{t(i)}}\;\sum_{l=1}^{L}\mathrm{CE}\!\big({\mathbf{x}}^{l},{\mathbf{x}}^{l}_{\varphi}(Z_{t(i)})\big)\bm{1}_{\{Z_{t(i)}^{l}={\mathbf{m}}\}}.$
Factorizing (B) for $L>1$ and following the results from above, we get
${\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ $=\mathbb{E}_{q(Z_{0}|{\mathbf{x}})}[-\log p_{\varphi}({\mathbf{x}}|Z_{0},{\mathbf{y}})]+\mathrm{KL}\!\big(q(Z_{1}|{\mathbf{x}})\,\|\,p_{\varphi}(Z_{1}|{\mathbf{y}})\big)$
@@
$)]\displaystyle\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}{,Z_{s(i)}}|{\mathbf{x}})}\Big[\sum_{l=1}^{L}\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z^{l}_{s(i)}}$$)\Big$]
$\quad{\;+\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Big[\sum_{l=1}^{L}\log\mathcal{Z}^{l}_{\varphi}(Z_{t(i)},\mathbf{y})\Big]}.$
@@
${\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ $=\mathbb{E}_{q(Z_{0}|{\mathbf{x}})}[-\log p_{\varphi}({\mathbf{x}}|Z_{0},{\mathbf{y}})]$
@@
$)]\displaystyle\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}{,Z_{s(i)}}|{\mathbf{x}})}\Big[\sum_{l=1}^{L}\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z^{l}_{s(i)}}$$)\Big$]
$\quad{\;+\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Big[\sum_{l=1}^{L}\log\mathcal{Z}^{l}_{\varphi}(Z_{t(i)},\mathbf{y})\Big]}.$
@@
${\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ $=\mathbb{E}_{q(Z_{0}|{\mathbf{x}})}[-\log p_{\varphi}({\mathbf{x}}|Z_{0})]$
@@
$)]\displaystyle\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}{,Z_{s(i)}}|{\mathbf{x}})}\Big[\sum_{l=1}^{L}\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z^{l}_{s(i)}}$$)\Big$]
$\quad{\;+\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Big[\sum_{l=1}^{L}\log\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})\Big]},$
which upon regrouping completes the proof of the statement. ∎
Implications. Theorem B.1 provides a principled upper bound on the negative log-posterior likelihood in discrete masked diffusion models.
  • Trainable upper bound. The pathwise upper bound ${\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ in (B.1) provides a principled training criterion for discrete diffusion posterior sampling.
@@
  • Data-consistency via tilt. The additive terms $\sum_{i}\mathbb{E}_{q(Z_{t(i)}{,Z_{s(i)}}|{\mathbf{x}})}[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))]$ encourage reverse transitions that produce intermediate predictions consistent with the measurement model, integrating task specific information at every denoising step.
@@
  • Compatibility with efficient parameterizations. Because (B.1) is written in terms of token-wise categoricals, it directly supports time-independent or lightweight parameterizations (e.g., shared denoisers), helping scalability to long sequences and high resolution.
Theorem B.2 (Test-time Anchored Posterior Sampling).
@@
${\mathcal{L}}_{\mathrm{APS}}({\mathbf{x}},{\mathbf{y}};\varphi)\coloneqq{\mathcal{L}}_{\mathrm{NELBO}}({\mathbf{x}};\theta)$ $+\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\!\Bigg[\frac{\alpha_{s(i)}-\alpha_{t(i)}}{1-\alpha_{t(i)}}\,\sum_{l=1}^{L}\log\frac{\langle{\mathbf{x}}^{l}_{\theta}(Z_{t(i)}),{\mathbf{x}}^{l}\rangle}{\langle{\mathbf{x}}^{l}_{\varphi}(Z_{t(i)}),{\mathbf{x}}^{l}\rangle}\,\bm{1}_{\{Z_{t(i)}={\mathbf{m}}\}}$
$)]\displaystyle\qquad\qquad\qquad\qquad-\mathbb{E}_{q({Z_{s(i)}}|Z_{t(i)},{\mathbf{x}})}\Big[\sum_{l=1}^{L}\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z^{l}_{s(i)}}$$)\Big$]
$\qquad\qquad\qquad\qquad{\;+\sum_{l=1}^{L}\log\mathcal{Z}^{l}_{\varphi}(Z_{t(i)},\mathbf{y})}\Bigg].$